Please solve using LaPlace partial fraction expansion the following equation (please write very neat, thank you):

d2y/dt2+4dy/dt+3y=0 y(0)=0 and y'(0)=12
Answer #1

Given Problem is dy +4 dy at + 3y = 0, 900)=0 dta y (0-12 data Apply Laplace Transform day { +44 +3493 – 2003 gely (8) -s960)On 12 = A (5+1) + B (5+3) 12 = (A+B) + CA+ 38) comparing coefficientwise AAB A+BB = 12 A = -6 =) BE6 - 2B -12. there for 6 Y(